Write the equation for the parabola in the form y=a(x-h)^2 +k with a y-intercept 10, x-intercept 2, and equation of axis of symmetry x-3=0
please help.

Answer by scott8148(6628)

from the axis of symmetry ___ h = 3

from the y-intercept (x = 0) ___ 10 = a(0 - 3)^2 + k ___ 10 = 9a + k

from the x-intercept (y = 0) ___ 0 = a(2 - 3)^2 + k ___ 0 = a + k

subtracting equations ___ 10 = 8a ___ 5/4 = a

substituting ___ 0 = (5/4) + k ___ -5/4 = k